OpenStudy (amenah8):

Find f(0.5), f(-6), and f(x+1) for the following functions

OpenStudy (amenah8):

\[f(x)-g(x)=x ^{2}+x-6\] and for g(x)=x+4

OpenStudy (triciaal):

substitute for x in the function and simplify example f(1/2) - g(1/2) = (1/2)^2 + (1/2) - 6 g(1/2) = (1/2) + 4 f(1/2)= (1/2)^2 +(1/2) - 6 + (1/2) + 4 = 1/4 + 5 - 6 = -1 + 1/4 = -3/4

OpenStudy (triciaal):

check calculations think I have an error but the method is OK

OpenStudy (amenah8):

thank you so much!! i understand it now

OpenStudy (triciaal):

f(1/2) = 1/4 + 1/2 - 6 + 9/2 1/4 + 2/4 -24/4+ 18/4 = -4/4 = -1

OpenStudy (triciaal):

g(-6) = -6 + 4 = -2 f(-6) = (-6)^2 + (-6) -6 + (-2) 36 -6 -6 -2 = 36 - 14 = 22

OpenStudy (triciaal):

g(x + 1) = (x + 1 ) + 4 = x + 5 f(x + 1 ) = (x +1)^2 + (x + 1) - 6 + (x + 5) x^2 + 2 x + 1 + x + 1 -6 + x + 5 = x^2 + 4 x + 1